Accessibility Assistance

Skip to Content




Information Technology
George A. Smathers Libraries
University of Florida
Suite 1250
2124 NE Waldo Rd
Gainesville, FL 32609

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
December 7, 2010

Present: Shruti Bagri, Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Winston Harris, Michael Jay, Tom Minton, Amy Polk, Patricia Ruwell, Josh Spurgin, Mark Sullivan, Evan Wack

Guest: Rachel Schipper
Announcements
  1. Rachel inquires about the status of the vaults in the old Library East IT Dept. Will notes the vaults are empty and the key returned to Rich Bennett. What about the old server room? There are still a number of items left to be surveyed in there. There are still some active servers as well. We are in the process of moving those to a CNS joint location. Hope to have the server equipment out by January 1st, but it depends, in part, on CNS. What is the status of the reel tapes stored in that room? Status is the same as it has been but Will can investigate and try to process out tapes that are no longer needed. We need to know what tapes can be disposed of and what tapes need to be retained. We know the Chiles tapes are to be retained (thought to be 8 or so tapes) and Joe Aufmuth's Map tapes are to be identified (it was thought that Joe was seeking external funding sources for conversion but there is uncertainty as to which tapes are involved). Rachel notes that they would like to repurpose the room. Tapes to be retained could be stored out at ALF- depending upon how many are to be retained. Will Chaney will check the email to see which tapes were identified for discard.
  2. Holiday party tomorrow in room 1A Library East. Rides available.
  3. Friday, December 17 - IT Staff luncheon. Discussion on where to dine. Decision: Merlion Singapore Cuisine on south 13th Street. Noon time.
Review Minutes of 11/30/10
  1. Software #3. It should be "Google Open ID (for secured login to the application)" and not Google Search.
Ongoing IT Updates

Hardware
  1. Software recall done. Debra started going through the software inventory. In some cases, not all the disks in a series have been returned. Hope to have a count of all inventoried software items soon.
  2. New Altiris server - broke. Logan to reinstall.
  3. Staff Win7 migration - planning now. Putting project together.
  4. MSL 107 technology proposal - Diane had questions about the equipment. We are waiting for her response. The quote seemed high.
  5. Rachel was inquiring about the videoconferencing needs for Education. She and Diane need an updated quote.
  6. Public migration - back burner. Did not receive the new Clean Slate build in time to test and get it out. Will shoot for spring - along with public laptop printing.
  7. Remedy - waiting for upgrade decision from UF Help Desk folk.
  8. SharePoint site - on hold until we have another tech hired.
  9. 1A air flow issue - still waiting on Facilities. All internal mics are muted, but they work externally.
  10. IT Specialist position closes tomorrow. There are a lot of applicants. Hope to start phone interviews next week and have someone hired before the break.
Software
  1. Statistics for Admin - number of public log-ons: 500,000 per year. Circulation laptops are not recorded. Stats are needed for budgeting justification. Website stats are also needed.
  2. FilmLog - on hold waiting for Preservation.
  3. Recruitment Study - Brian is happy with the progress. He has requested some changes. Need report.
  4. Document Disposition database - demo to GovDocs this week. Next week, will demo to GPO talks in Washington, DC. Created a LibGuide with instructions. Aiming for a January launch in Florida (40 institutions).
  5. Cathy Martyniak is now requesting a Journals/Storage Disposition database.
  6. ILL server - Chris is working with Chris Youngblood to switch to the new LIB-ILL server and off NT18.
  7. Macros - Michael has a handful of active ones awaiting user input. He is experiencing a problem getting response from users. Give Will a list of the non-responsive users. It should not linger too long or someone should speak with the department chair.
  8. Kuali OLE - Maureen Kelly is out today. They need a lead architect. Programmers will be starting sprints. Kuali has a financial product out that we would like to work with OLE. There are currently six Kuali products in production.
Network/Server Administration
  1. ALF network/speed issues - CNS has been tweaking it. No major slowness reported. It's always slow into campus. CNS can't find a problem. Usage grafts do not show problems. They'll keep working on it. Hope to get fiber optics out there eventually. When they have more staff out there, it will be a bigger issue. Need a better estimate for the fiber. Will ask about the status.
  2. ALF - interim storage area - perhaps the old Lowes building on Waldo Road.
  3. SharePoint - going well. Circ Manual site created. Aleph Pro - need a better idea of what is needed. It's in web page form now. May need wiki on SharePoint. Will set up structure and access. Have a meeting tomorrow with Trey Shelton and Raimonda regarding the Acquisitions site. Cataloging is moving forward. MSL is in use. Administration is underway. Logan to provide instruction to Admin on site usage.
  4. Server room - NT18 - waiting for ILL to give the okay to switch to new server and power down the old.
  5. GIS server - finished last week. Joe says it all went well. Moving files off NT13. West security - waiting for the cost of the co-location with CNS.
  6. Revamp on WordPress server - added social networking features, updated themes. Changed commenting system. Changes should help to minimize spam. It will now be more useful and interactive for students and staff. Upgraded to WordPress 3.02.
  7. CNS virtualization is complete. Move to CNS - will need to take down EPO. It can't be down for two days. Will need to be moved to a temporary location in the interim.
ILS
  1. Cynthia is plugging away.
Web Projects
  1. Acquisitions reorg - Steve Carrico asked for the ticket to be closed.
  2. Second Acquisitions project: ARL stats. Waiting for Excel tables to put up. Includes Law, HSCL, etc.
  3. Proxy server issues (3 parts): failover (done), Jason project with transparent URLs, and the 14-digit number in Shibboleth. The only way around the 14-digits is to use VPN. Decision on the proxy is problematic. The trend seems to be going towards VPN. Transparent URLs - would need test proxy servers.
  4. Two more surveys put up this week.
  5. Jan pointed out that the extended hours on the web page were wrong. The app shows the Library as closed. Will look into it. Marlo is entering the hours. Check on hours for spring as well.